OpenStudy (anonymous):

whats x? 7(x-5) + 14=7

OpenStudy (anonymous):

x=4

OpenStudy (lalaly):

7x-35+14=7 7x-21=7 7x=28 x=4

OpenStudy (anonymous):

7(x-5)+14=7 subtract 14 from both sides 7(x-5)=-7 divide by 7 on both sides (x-5)=-1 add 5 to both sides x=4

OpenStudy (anonymous):

7x-35+14=7 7x-21=7 7x=28 x=4

OpenStudy (anonymous):

x = 4